1st Question:-

What is the contraction of the Na and K if we take 1ml from Na (10ppm) and 1ml from K (10ppm) and dissolved them together with dist. water to be 10 ml mixed soln.?





2nd Question:-

What is the contraction of the Na and K if we take 1ml from a mixture contains Na (10ppm) and K(10ppm) and dissolved it with dist. water to be 10 ml mixed soln.?
1
Expert's answer
2013-06-04T11:24:29-0400
1) 1mL = 0.01 mg Na, 10 ppm Nа
1 mL = 0.01 mg К, 10 ppm К
To calculate the concentration we can use the formula
C(Na) = n(Na) / V solution
C(K) = n(K) / V solution
To calculate the amount of substance of Na and K we can use the formula
n(Na) = m/M
n(K) = m/M
n(Na) = 0.0001 g / 23 g/mol = 4.35e-6 mol
n(K) = 0.0001 g / 39 g/mol = 2.56e-6 mol
So the concentration will be:
C(Na) = 4.35e-6 mol / 0.01 L = 0.000435 mol/L
C(K) = 2.56e-6 mol / 0.01 L = 0.000265 mol/L
Answer: 0.000435 mol/L, 0.000265 mol/L
2) 1 mL = 0.005 mg Na + 0.005 mg K
To calculate the concentration we can use the formula
C(Na) = n(Na) / V solution
C(K) = n(K) / V solutionto calculate the amount of substance of Na and K we can use the formulan(Na) = m/M
n(K) = m/M
n(Na) = 0.00005 g / 23 g/mol = 2.17e-6 mol
n(K) = 0.00005 g / 39 g/mol = 1.28e-6 mol
So the concentration will be:
C(Na) = 2.17e-6 mol / 0.01 L = 0.000217 mol/L
C(K) = 1.28e-6 mol / 0.01 L = 0.000128mol//L
Answer: 0.000217mol/L, 0.000128mol/L